The area under one arch of the sine curve revolves about the x-axis. Find the volume generated.


1
Expert's answer
2021-08-03T12:36:26-0400

Solution;

Use the disc method;

"V=π\int_a^b(f(x))^2dx"

Given;

f(x)=sin(x)

For one arch of the sine curve function,the limits varies from 0 to π

a=0

b=π

Substitute in the formula;

"V=π\int_0^π(sin(x))^2dx"

"V=π\int_0^2(sin^2(x)dx"

By the trigonometric identity;

cos(2x)=1-2sin2(x)

Gives; "sin^2(x)=\frac{1-cos(2x)}{2}"

By substitution;

"V=π\int_0^π(\frac{1-cos(2x)}{2})dx"

"V=\fracπ2[\int_0^π(1-cos(2x))dx"

"V=\fracπ2[\int_0^π1dx-\int_0^πcos(2x)dx"

"V=\fracπ2[x]_0^π-\fracπ2[\frac{sin(2x)}{2}]_0^π"

"V=\fracπ2[π]-\fracπ4[sin(2π)-sin(0)]"

"V=\frac{π^2}{2}-\fracπ4(0)"

"V=\frac{π^2}{2}"

Answer;

"\frac{π^2}{2}" cubic units.